# Tales of the 2025 Los Angeles Fire: Hotwash for Public Health Concerns in Reddit via LLM-Enhanced Topic Modeling

### v2

@@ 3.3.2 Comment-level Modeling
Step 1: Sentence embedding. We utilized the all-mpnet-base-v2 model from the SentenceTransformers library to encode each comment into dense vector representations. This general-purpose model demonstrated the best performance at sentence level. Ashqar and Mutlu (2023).
@@
[…]
@@ 5.1 Revealing Public Health Concerns
Back to the effective datasets, approximately half of all identified instances are directly related to public health and safety, intersecting with other SA categories, such as fire operations, emergency resources and loss and damage. Sepecifically, we observe multiple latent topics in the public health and safety category. These include concerns about environmental health, particularly regarding air and water quality; occupational health, with a focus on firefighters, journalists, and emergency responders; and one health, debating on how the use of seawater could nagetively impact inland vegetation. Furthermore, several discussions highlight disparities in access to real-time medical support and protective resources. These findings align with prior research emphasizing the adverse health outcomes of wildfire smoke exposure, particularly among vulnerable populations Rappold et al. (2017); Zhang et al. (2025); Lei et al. (2024).
@@ 5.2 Grief and Mental Health Risks Detection
The crisis narrative category captures the complex emotions and reactions of individuals in such an unexpected natural diaster. Notably, our temporal analysis suggests an absence of a preparedness phase because there is no instance on Reddit before January 7. In general, there should be 3 stages, preparedness, response and recovery, however, there is no discussion before January 7. The gap between anticipated and actual outcomes can lead to cognitive dissonance Dwivedi et al. (2018), as individuals struggle to reconcile their proactive measures with the devastation. Consequently, learned helplessness may erode confidence in future resilience efforts within affected communities Landry et al. (2018); Ramachandran et al. (2021). Similarly, we detect strong signal of grief emotion and mental health risks in the Reddit data. While the physical spread of the fire lasted only a few days, emotional and psychological recovery may take far longer as other primary mental health studies discovered in many wildfire events around the world Wettstein and Vaidyanathan (2024); To et al. (2021).
@@ 6 Conclusion
{mdframed}
[backgroundcolor=gray!10, linewidth=1pt, leftmargin=0pt, innerleftmargin=10pt, innerrightmargin=10pt] Topic 3: Air Quality and Health Concerns Post-Wildfires SA: Public health and safety, Emergency resources, Recovery, Loss and damage CN: Victim, Blame, Renewal Grief Signal: Yes  Mental Health Risk: Yes Topic 24: Evacuation Strategies: Cars vs. Buses in Emergencies SA: Public health and safety, Emergency resources, Loss and damage CN: Blame, Victim Grief Signal: No  Mental Health Risk: No
@@
The data and code will be made publicly available after the journal submission.

### v3

@@ 3.3.2 Comment-level Modeling
Step 1: Sentence embedding. We utilized the all-mpnet-base-v2 model from the SentenceTransformers library to encode each comment into dense vector representations. This general-purpose model demonstrated the best performance at sentence level. ashqar2023comparative.
@@
In steps 2 and 3, we tuned hyperparameter, shown in table 4 and supplementary figure S2, across n_neighbors $\in$ {15, 20, 25, 30}, min_dist $\in$ {0.0, 0.01}, and min_cluster_size $\in$ {50, 100, 150, 200, 250, 300, 350, 400}, while setting min_samples to half of the cluster size. For each configuration, we trained a separate BERTopic model and evaluated its performance using the best-performing coherence metric $C_{v}$ roder2015exploring with the CoherenceModel from the Gensim library. The final model, configured with n_neighbors = 30, min_dist = 0.01, and min_cluster_size = 300, achieved the best balance between topic coherence (0.628), interpretability, and semantic diversity. The optimal number of latent topics is 30, we further mapped the 30 latent topics to the pre-defined hierarchical category forwork, the mapping examples are shown in Supplementary Information S3.
@@ 5.1 Revealing Public Health Concerns
Back to the effective datasets, approximately half of all identified instances are directly related to public health and safety, intersecting with other SA categories, such as fire operations, emergency resources and loss and damage. Sepecifically, we observe multiple latent topics in the public health and safety category. These include concerns about environmental health, particularly regarding air and water quality; occupational health, with a focus on firefighters, journalists, and emergency responders; and one health, debating on how the use of seawater could nagetively impact inland vegetation. Furthermore, several discussions highlight disparities in access to real-time medical support and protective resources. These findings align with prior research emphasizing the adverse health outcomes of wildfire smoke exposure, particularly among vulnerable populations rappold2017community; zhang2025respiratory; lei2024wildfire.
@@ 5.2 Grief and Mental Health Risks Detection
The crisis narrative category captures the complex emotions and reactions of individuals in such an unexpected natural diaster. Notably, our temporal analysis suggests an absence of a preparedness phase because there is no instance on Reddit before January 7. In general, there should be 3 stages, preparedness, response and recovery, however, there is no discussion before January 7. The gap between anticipated and actual outcomes can lead to cognitive dissonance dwivedi2018involvement, as individuals struggle to reconcile their proactive measures with the devastation. Consequently, learned helplessness may erode confidence in future resilience efforts within affected communities landry2018learned; ramachandran2021learned. Similarly, we detect strong signal of grief emotion and mental health risks in the Reddit data. While the physical spread of the fire lasted only a few days, emotional and psychological recovery may take far longer as other primary mental health studies discovered in many wildfire events around the world wettstein2024psychotropic; to2021impact.
@@ 6 Conclusion
Topic 3: Air Quality and Health Concerns Post-Wildfires SA: Public health and safety, Emergency resources, Recovery, Loss and damage CN: Victim, Blame, Renewal Grief Signal: Yes  Mental Health Risk: Yes Topic 24: Evacuation Strategies: Cars vs. Buses in Emergencies SA: Public health and safety, Emergency resources, Loss and damage CN: Blame, Victim Grief Signal: No  Mental Health Risk: No
@@
All code used in this study is publicly available on GitHub ( https://github.com/MoonSulong/2025WildfireCA). The de-identified datasets generated and analyzed during the study (text only; all images were removed to protect user privacy) are hosted on Hugging Face (https://huggingface.co/datasets/Dragmoon/2025CalifoniaWildfire).